# Proximity Effects Between the Graphene Quasicrystal and Magic-Angle Twisted Bilayer Graphene

### v2

@@ Abstract
We report a numerical study of graphene heterostructures comprised of three individual layers twisted by either the magic angle of 1.1∘, or , corresponding to the graphene quasicrystal. The heterostack is modeled using realistic structural and tight-binding models, while transport properties are calculated in both clean and disordered systems containing up to 8 million atoms. The weak interaction between different layers allows us to scrutinize the electronic mixing of flat bands and quasicrystalline states, which are altered differently in the low- and high-energy regimes and provide a new type of hybrid physics to be explored.
Introduction. The fascinating discovery of the superconducting and Mott insulating phases in magic-angle twisted bilayer graphene (MATBLG) has generated a huge interest in deepening our understanding of the physics of flat-band systems, where a dominant Coulomb interaction triggers a wealth of emerging many-body phenomena [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. In parallel, a “graphene quasicrystal” for a twist angle of 30∘ between two graphene layers has been observed [18], and later grown deterministically via chemical vapor deposition [19], eliminating the need for artificial assembly. In the decades after their first experimental observation [20], the realization of quasicrystalline structures has captivated material scientists for their quasiperiodic tiling with self-similar properties [21, 22, 23, 24]. Very recently, other generalizations of Moiré quasicrystals have been also discovered [25, 26].
While the exotic physics of MATBLG is dominated by low-energy flat bands, the peculiar electronic states in graphene quasicrystals are to date mainly found at high energies, and the low-energy region seems to be well described by the physics of the individual graphene layers. The combination of both types of stacking could thus generate a strange mixture of “two different worlds,” which are separately still under intense investigation.
@@
In this Letter, we overcome these difficulties by implementing a realistic tight-binding model which describes the electronic properties of graphene layers stacked at both the magic angle and the quasicrystalline approximant twist angles. In combination with an efficient linear-scaling approach [29], we simulate the electronic transport properties of MATBLG/graphene quasicrystal trilayer stacks containing up to 8 million atoms, thus allowing for the first time the study of the respective modifications of flat bands and quasicrystalline-like states driven by proximity effects.
In the absence of disorder, the flat bands of MATBLG are found to be unaffected by the quasicrystalline states. This contrasts with the high-energy quasicrystalline states, which are suppressed by the interaction with MATBLG. However, when introducing disorder we see the opposite effect. In this case, in the low-energy flat bands we observe a reversal of the recently discovered disorder-induced delocalization in MATBLG [30]. Meanwhile, in the high-energy quasicrystalline states, the transport properties are unaffected by the presence of the magic-angle graphene layer.
Structural optimization and tight-binding model. The atomic structures of graphene superlattices at the magic angle and at the quasicrystal approximant angles were relaxed using molecular dynamics simulations with classical potentials [31, 32]. Intralayer forces were computed using the optimized Tersoff potentials [33], whereas interlayer forces were modeled using the Kolmogorov–Crespi potentials [34, 35]. The atomic structures were optimized until all force components were smaller than 0.5 meV/Å. To model quasicrystalline graphene, we consider two approximant angles close to 30∘, namely 31∘ and 29.8∘ [27]. These approximants share similarities with the fully aperiodic quasicrystal, such as the formation of high energy states whose real space projections exhibit dodecagonal symmetry [18].
@@
Calculation of electronic properties. We use an efficient linear-scaling quantum transport methodology [29] to study the electronic properties of the MATBLG/quasicrystal trilayer stacks. For transport, the main quantity we calculate is the mean square displacement of an initial state , given by
@@
The above results have shown that, in the clean limit, adding a -rotated graphene layer to MATBLG has little impact on the low-energy flat-band states. We now see if this extends to transport in the presence of disorder. We consider Anderson disorder values, and , where we see an anomalous disorder-induced delocalization in MATBLG, as discovered in Ref. 30. We reproduce these results in the inset of Fig. 4, where we show the time evolution of the diffusion coefficient of MATBLG in the middle of the flat band (). Here we see that when increasing the disorder strength, we see an anomalous increase in at long times.
This situation is strongly altered in the trilayer case, as shown in the main panel of Fig. 4. Here we see that at long times, increasing the disorder strength always reduces the diffusion coefficient. The anomalous disorder-induced delocalization seen in MATBLG thus appears to be suppressed when adding a quasicrystalline graphene layer to the stack, resulting in “standard” transport behavior. This occurs despite the strong resilience of the state projection on the MATBLG layers in the clean case.
High-energy regime and quasicrystalline states. We now focus on the electronic structure and transport in the high-energy regime. Here we obtain a favorable comparison to prior results on graphene quasicrystalline approximants [40, 27], which have demonstrated the presence of resonant peaks in the DOS at energies between and eV, including the so-called -peak shown as the vertical dotted line in Fig. 5. These peaks correspond to localized states lacking spatial periodicity and displaying quasiperiodic orientational patterns. The top panel of Fig. 5 shows that adding an extra -rotated layer on top of the quasicrystalline stack leads to strong suppression of the quasicrystalline fingerprints in the DOS.
@@
Finally, we focus on transport in the high-energy regime, where the formation of resonant quasicrystalline states has been observed experimentally and is well reproduced theoretically. At those resonance energies, the real-space extension of the electronic states indicates the emergence of dodecagonal ordering (incompatible with translational invariance), which could point towards unconventional electronic features in the limit of a totally aperiodic structure, as known for other quasicrystalline structures [41, 42, 43, 44, 45, 46]. Our simulations, however, do not reveal any unusual transport behavior in the graphene quasicrystalline approximant. As shown in Fig. 6, Anderson disorder leads to standard diffusive behavior, with stronger disorder leading to slower diffusion and a reduced mean free path (shown in the inset). We also find that, in contrast to the DOS and the Fermi velocity in the clean case, adding a -rotated layer has minimal impact on transport in the presence of disorder.
Summary and conclusions. We have shown that interfacing MATBLG with graphene quasicrystal approximants leads to a variety of proximity effects. While the low-energy flat-band states appear to be insensitive to the addition of a -rotated layer in the clean case, the high-energy quasicrystalline states are suppressed by the presence of a magic-angle layer. Meanwhile, adding disorder reveals the opposite trend. In the low-energy regime, the proximity effect suppresses the anomalous disorder-induced delocalization discovered in the flat bands of MATBLG [30]. However, in the high-energy regime, disorder does not lead to any transport anomalies in the quasicrystal states, independent of the presence of the magic-angle layer. This contrasts with the physics of intermetallic alloys-based quasicrystals, where disorder-induced delocalization remains a long-standing conundrum [47, 41, 48]. Looking ahead, it would be instructive to deepen the study of disorder effects in light of recent developments in other quasicrystalline systems [42, 43, 44, 45, 46].
@@ Acknowledgements.
ICN2 is funded by the CERCA Programme/Generalitat de Catalunya and supported by the Severo Ochoa Centres of Excellence programme, Grant CEX2021-001214-S, funded by MCIN/AEI/10.13039.501100011033. This work is also supported by MICIN with European funds‐NextGenerationEU (PRTR‐C17.I1) and by 2021 SGR 00997, funded by Generalitat de Catalunya. V.-H.N. and J.-C.C. acknowledge financial support from the Fédération Wallonie-Bruxelles through the ARC Grant “DREAMS” (No. 21/26-116), from the EOS project “CONNECT” (No. 40007563), from the EU Pathfinder project “FLATS” (No. 101099139), and from the Belgium F.R.S.-FNRS through the research project (No. T.029.22F). Simulations were performed at the Center for Nanoscale Materials, a U.S. Department of Energy Office of Science User Facility, supported by the U.S. DOE, Office of Basic Energy Sciences, under Contract No. 83336. Computational resources were also provided by the supercomputing facilities of UCLouvain (CISM) and the Consortium des Équipements de Calcul Intensif (CÉCI), funded by the F.R.S.-FNRS under Grant No. 2.5020.11 and by the Walloon Region. We acknowledge grant PID2022-138283NB-I00 funded by MICIU/AEI/ 10.13039/501100011033 and by “ERDF/EU”.References
@@
- Fan et al. [2021] Z. Fan, J. H. Garcia, A. W. Cummings, J. E. Barrios-Vargas, M. Panhans, A. Harju, F. Ortmann, and S. Roche, Linear scaling quantum transport methodologies, Phys. Rep. 903, 1 (2021), linear scaling quantum transport methodologies.
- [30] P. A. Guerrero, V.-H. Nguyen, J. M. Romeral, A. W. Cummings, J.-H. Garcia, J.-C. Charlier, and S. Roche, Disorder-induced delocalization in magic-angle twisted bilayer graphene, arXiv:2401.08265 .
@@
- [44] T. Peng, Y.-C. Xiong, C.-B. Hua, Z.-R. Liu, X. Zhu, W. Cao, F. Lv, Y. Hou, B. Zhou, Z. Wang, and R. Xiong, Structural disorder-induced topological phase transitions in quasicrystals, arXiv:2403.04338 .

### v3

@@ Abstract
We present a numerical study of three-layer graphene heterostructures in which the layers are twisted by the magic angle (1.1∘) or by to form a graphene quasicrystal. The heterostacks are described using realistic structural relaxations and tight-binding Hamiltonians, and their transport properties are computed for both pristine and disordered systems containing up to 8 million atoms. Owing to the weak interlayer coupling, we resolve the hybridization between magic-angle flat bands and quasicrystalline states, which are modified in distinct ways across low- and high-energy windows, revealing a new hybrid electronic regime to explore.
Introduction. The fascinating discovery of the superconducting and Mott insulating phases in magic-angle twisted bilayer graphene (MATBLG) has generated a huge interest in deepening our understanding of the physics of flat-band systems, where a dominant Coulomb interaction triggers a wealth of emerging many-body phenomena [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. In parallel, a “graphene quasicrystal” for a twist angle of 30∘ between two graphene layers has been observed [18], and later grown deterministically via chemical vapor deposition [19], eliminating the need for artificial assembly. In the decades after their first experimental observation [20], the realization of quasicrystalline structures has captivated material scientists for their quasiperiodic tiling with self-similar properties [21, 22, 23, 24, 25]. Very recently, other generalizations of Moiré quasicrystals have been also discovered [26, 27].
MATBLG is dominated by low-energy flat bands, whereas graphene quasicrystals display their most distinctive states at high energies, with low-energy behavior resembling that of isolated graphene layers. Their combination could create a nontrivial hybrid of these two regimes, both still actively explored.
@@
Here we overcome these difficulties by implementing a realistic tight-binding model which describes the electronic properties of graphene layers stacked at both the magic angle and the quasicrystalline approximant twist angles. In combination with an efficient linear-scaling approach [30], we simulate the electronic transport properties of the graphene quasicrystal and the MATBLG/graphene quasicrystal trilayer stacks containing up to 8 million atoms, thus allowing for the first time the study of the respective modifications of flat bands and quasicrystalline-like states driven by proximity effects.
In the absence of disorder, the flat bands of MATBLG are found to be unaffected by the quasicrystalline states. This contrasts with the high-energy quasicrystalline states, which are severely suppressed by the interaction with MATBLG. However, this changes when introducing disorder. In this case, in the low-energy flat bands we observe a reversal of the recently discovered disorder-induced delocalization in MATBLG [31]. Meanwhile, in the high-energy quasicrystalline states, anomalous (sub-ballistic) transport properties (typical to pristine quasiperiodic ordering) are always suppressed by disorder, independent of the presence of the magic-angle graphene layer. Overall, our results indicate that a third graphene layer is generally detrimental to observing the anomalous transport features of both the low-energy flat bands in MATBLG and the high-energy quasicrystal states in disorder-free -BLG.
Structural optimization and tight-binding model. The atomic structures of graphene superlattices at the magic angle and at the quasicrystal approximant angles were relaxed using molecular dynamics simulations with classical potentials [32, 33]. Intralayer forces were computed using the optimized Tersoff potentials [34], whereas interlayer forces were modeled using the Kolmogorov–Crespi potentials [35, 36]. The atomic structures were optimized until all force components were smaller than 0.5 meV/Å. To model quasicrystalline graphene, we consider two approximant angles close to 30∘, namely 31∘ and 29.8∘ [28]. These approximants share similarities with the fully aperiodic quasicrystal, such as the formation of high-energy states whose real space projections exhibit dodecagonal symmetry [18].
@@
Calculation of electronic properties. We use an efficient linear-scaling quantum transport methodology [30] to study the electronic properties of the MATBLG/quasicrystal trilayer stacks and the two involved bilayers. For transport, the main quantity we calculate is the mean square displacement of an initial state , given by
@@
The above results have shown that, in the clean limit, adding a -rotated graphene layer to MATBLG has little impact on the low-energy flat-band states. We now see if this extends to transport in the presence of disorder. We consider Anderson disorder values, and , where we see an anomalous disorder-induced delocalization in MATBLG, as discovered in Ref. 31. We reproduce these results in the inset of Fig. 4, where we show the time evolution of the diffusion coefficient of MATBLG in the middle of the flat band (). Here we see that increasing the disorder strength leads to an anomalous increase in at long times.
This situation is strongly altered in the trilayer case, as shown in the main panel of Fig. 4. Here we see that at long times, increasing the disorder strength always reduces the diffusion coefficient. The anomalous disorder-induced delocalization seen in MATBLG is thus suppressed when a quasicrystalline graphene layer is added to the stack, resulting in “standard” transport behavior. This occurs despite the strong resilience of the state projection on the MATBLG layers in the clean case. The presence of disorder thus appears to enhance the coupling between the layers and wash out the flat bands by enabling scattering between the flat-band states and higher-energy states in the third layer.
High-energy regime and quasicrystalline states. We now focus on the electronic structure and transport in the high-energy regime. Here we obtain a favorable comparison to prior results on graphene quasicrystalline approximants [41, 28, 42], which have demonstrated the presence of resonant peaks in the DOS at energies between and eV, including the so-called -peak shown as the vertical dotted line in Fig. 5. These peaks correspond to localized states lacking spatial periodicity and displaying quasiperiodic orientational patterns. The top panel of Fig. 5 shows that adding an extra -rotated layer on top of the quasicrystalline stack leads to strong suppression of the quasicrystalline fingerprints in the DOS.
@@
The LDOS of the pristine quasicrystal bilayer, shown in Fig. 6 (top), indicates the emergence of the fractal localization pattern, consistent with prior results [18, 28, 42]. In contrast, the LDOS of the equivalent atomic coordinates in the trilayer does not exhibit any anomalous localization features, indicating that the addition of an extra layer destroys the quasicrystalline states and their related features.
@@
Finally, we focus on transport in the high-energy regime, where the formation of resonant quasicrystalline states has been observed experimentally and is well reproduced theoretically. At those resonance energies, the real-space extension of the electronic states indicates the emergence of dodecagonal ordering (incompatible with translational invariance), which could point towards unconventional electronic features in the limit of a totally aperiodic structure, as known for other quasicrystalline structures [22, 43, 44, 45, 46, 47, 48]. We can spot signals of this unconventional transport in the short time behavior shown in the inset of Fig. 7 (top panel). Here, the quantum dynamics in the quasicrystalline approximant (indicated by the blue line) exhibit the typical sub-ballistic transport even in the absence of disorder, as observed in pure quasiperiodic systems [22, 49, 50, 51, 52, 53]. However, adding the third -layer restores ballistic transport in the absence of disorder (black line), consistent with the suppression of quasiperiodic-induced anomalous localization effects indicated in Figs. 5 and 6.
Additionally, as shown in the top panel of Fig. 7, Anderson disorder leads to standard diffusive behavior for both the bilayer quasicrystal and the trilayer, with a stronger disorder leading to slower diffusion and a reduced mean free path. The bottom panel shows the LDOS of the approximant with a superimposed Anderson disorder of strength . The fractal localization pattern present in the pristine case (top panel of Fig. 6) is clearly destroyed. These results, as well as those of the clean case, suggest that the graphene quasicrystal states are poorly resilient against local perturbation, whether arising from short-range electrostatic disorder or from coupling to another layer.
Summary and discussion. We have shown that interfacing MATBLG with graphene quasicrystal approximants leads to a variety of proximity effects. While the low-energy flat-band states appear to be insensitive to the addition of a -rotated layer in the clean case, the high-energy quasicrystalline states are suppressed by the presence of a magic-angle layer. Meanwhile, adding disorder reveals a couple different features. In the trilayer, the presence of disorder always leads to standard diffusive behavior, both in the low-energy flat bands and in the high-energy quasicrystalline states. Disorder also washes out the quasicrystalline states in the bilayer system, while an anomalous increase of mean free path with disorder strength is observed in the flat bands of MATBLG [31].
These results have a couple implications for heterostructures involving flat-band or quasicrystal states in twisted BLG. In the case of MATBLG, Figs. 2 and 3 suggest that a -rotated graphene layer can serve as a protective capping layer to protect flat-band states from extrinsic disorder or adsorbates. However, this holds only in the clean case. Once disorder is introduced, the anomalous features of the flat-band states are washed out, indicating that such a capping layer is ultimately more detrimental than beneficial. A semiconducting or insulating capping layer may be more useful for this purpose, as a large band gap may prevent disorder-induced mixing of the flat bands with higher-energy states of the capping layer.
In the quasicrystalline case, we find that the graphene quasicrystal state is destroyed both by disorder and by interaction with a third graphene layer, suggesting that it is highly fragile to local perturbations. This contrasts with the physics of intermetallic alloy-based quasicrystals, where disorder-induced delocalization remains a long-standing conundrum [49, 22, 50]. Here a difference may lie in the short- vs. long-range nature of the disorder; as the quasicrystal state exhibits localized features on the scale of the graphene lattice constant [28], it may be more susceptible to the short-range disorder we consider here. Meanwhile, long-range disorder, such as that arising from electron-hole puddles [54], may preserve the anomalous transport features of the quasicrystalline state.
A similar, but opposite, argument may be made in the case of MATBLG. Here the flat band states have a periodicity of nm [31], and short-range disorder may therefore average out over this length scale and be less effective. Meanwhile, electron-hole puddles, whose length scale is similar to the periodicity of the magic-angle Moiré [54], may be more effective in suppressing the anomalous transport features of the flat bands.
To summarize, our simulations indicate that a third graphene layer is generally detrimental to observing the anomalous transport features of both the low-energy flat bands in MATBLG and the high-energy quasicrystal states in -BLG. Looking ahead, our results open promising avenues for exploring these systems, particularly their response to the diverse types of defects and disorder commonly encountered in bilayer graphene, as well as the extension of proximity effects to other types of magnetic or strong spin-orbit coupling materials.
@@ Acknowledgements.
ICN2 is funded by the CERCA Programme/Generalitat de Catalunya and supported by the Severo Ochoa Centres of Excellence programme, Grant CEX2021-001214-S, funded by MCIN/AEI/10.13039.501100011033. This work is also supported by MICIN with European funds‐NextGenerationEU (PRTR‐C17.I1) and by 2021 SGR 00997, funded by Generalitat de Catalunya. V.-H.N. and J.-C.C. acknowledge financial support from the Fédération Wallonie-Bruxelles through the ARC Grant “DREAMS” (No. 21/26-116), from the EOS project “CONNECT” (No. 40007563), from the EU Pathfinder project “FLATS” (No. 101099139), and from the Belgium F.R.S.-FNRS through the research project (No. T.029.22F). Simulations were performed at the Center for Nanoscale Materials, a U.S. Department of Energy Office of Science User Facility, supported by the U.S. DOE, Office of Basic Energy Sciences, under Contract No. 83336 and 83864. Computational resources were also provided by the supercomputing facilities of UCLouvain (CISM) and the Consortium des Équipements de Calcul Intensif (CÉCI), funded by the F.R.S.-FNRS under Grant No. 2.5020.11 and by the Walloon Region.References
@@
- Fan et al. [2021] Z. Fan, J. H. Garcia, A. W. Cummings, J. E. Barrios-Vargas, M. Panhans, A. Harju, F. Ortmann, and S. Roche, Linear scaling quantum transport methodologies, Phys. Rep. 903, 1 (2021).
- Guerrero et al. [2025] P. A. Guerrero, V.-H. Nguyen, J. M. Romeral, A. W. Cummings, J.-H. Garcia, J.-C. Charlier, and S. Roche, Disorder-induced delocalization in magic-angle twisted bilayer graphene, Phys. Rev. Lett. 134, 126301 (2025).
@@
- Vidarte and Lewenkopf [2024] K. J. U. Vidarte and C. Lewenkopf, Quasicrystalline 30° twisted bilayer graphene: fractal patterns and electronic localization properties, Front. Carbon 3, 10.3389/frcrb.2024.1496179 (2024).
@@
- Peng et al. [2024] T. Peng, Y.-C. Xiong, C.-B. Hua, Z.-R. Liu, X. Zhu, W. Cao, F. Lv, Y. Hou, B. Zhou, Z. Wang, and R. Xiong, Structural disorder-induced topological phase transitions in quasicrystals, Phys. Rev. B 109, 195301 (2024).
@@
- [48] H. Roche-Carrasco, J. Schirmann, A. Mordret, and A. G. Grushin, A family of aperiodic tilings with tunable quantum geometric tensor, arXiv:2505.13304 .
@@
- Triozon et al. [2004] F. Triozon, S. Roche, A. Rubio, and D. Mayou, Electrical transport in carbon nanotubes: Role of disorder and helical symmetries, Phys. Rev. B 69, 121410 (2004).
- Triozon et al. [2002] F. Triozon, J. Vidal, R. Mosseri, and D. Mayou, Quantum dynamics in two- and three-dimensional quasiperiodic tilings, Phys. Rev. B 65, 220202 (2002).
- Trambly de Laissardière et al. [2026] G. Trambly de Laissardière, S. Venkateswarlu, A. Misssaoui, G. Jemaï, K. Chika, J. Vahedi, O. Faizy Namarvar, J.-P. Julien, A. Honecker, L. Magaud, J. Jemaa Khabthani, and D. Mayou, Electronic structure and transport in materials with flat bands: 2d materials and quasicrystals, Physica E 175, 116362 (2026).
- Xue et al. [2011] J. Xue, J. Sanchez-Yamagishi, D. Bulmash, P. Jacquod, A. Deshpande, K. Watanabe, T. Taniguchi, P. Jarillo-Herrero, and B. J. LeRoy, Scanning tunnelling microscopy and spectroscopy of ultra-flat graphene on hexagonal boron nitride, Nat. Mater. 10, 282 (2011).